# Scaling relations for heat and momentum transport in sheared Rayleigh–Bénard convection

### v1

@@ 3.3 Large scale circulation
Figure 6: Visualisation of $\boldsymbol{\tau_{w}^{\prime}}$ for the CRB system with $Ra=10^{7}$ and $Pr=1$. Plots (c), (e), and (g) are in the buoyancy-dominated regime and plot (i) is in the shear-dominated regime. The colour bars indicate the magnitude of $\boldsymbol{\tau_{w}^{\prime}}$ while the black arrows in the magnified panels (b), (d), (f), (h) and (j) indicate the direction of $\boldsymbol{\tau_{w}^{\prime}}$. The black lines in (c-h) indicate the slope of $\bigl{(}Re_{w}$/$Re_{R}\bigr{)}^{-1}$.
@@
Figure 7: The probability distribution $\phi(\alpha)$ of the flow orientation angle $\alpha$ for all heights $0\leq z/\lambda_{\theta}\leq Nu$ for $Ra=10^{7}$, $Pr=1$, $\Gamma_{x}=48$ and $\Gamma_{y}=24$ in (a) RB flow (i.e., with $Re_{S}=0)$, (b,d) buoyancy-dominated regime with $Re_{S}/Re_{R}\approx 2.11$, and (c,e) shear-dominated regime with $Re_{S}/Re_{R}\geq 10$. Plots (b,c) are for the CRB system while the plots (d,e) are for the PRB system. The black dashed lines indicate the height of the thermal boundary layer and the colour bar indicates the magnitude of the probability. The probability distribution $\phi(\alpha)$ at the thermal boundary layer height plotted against angle $\alpha$ for various $Ra$ and $Pr$ in (f) RB flow, (g) buoyancy-dominated regime, and (h) shear-dominated regime. The solid lines are the CRB system with $\Gamma_{x}=48$ and $\Gamma_{y}=24$, the dashed lines are for the CRB system with $\Gamma_{x}=8$ and $\Gamma_{y}=4$, and the dotted lines are for the PRB system with $\Gamma_{x}=8$ and $\Gamma_{y}=4$.
@@ 3.4 Dissipation rates
Figure 8: (a) Boundary layer contribution to the global kinetic dissipation rate $\epsilon_{u,BL}$ and (b) the boundary layer contribution to the global thermal dissipation rate $\epsilon_{\theta,BL}$. The circle markers are for the CRB system, while plus markers are for the PRB system.
We now investigate the bulk and boundary layer contributions to the global kinetic and thermal dissipation rates described in §2.1. For the range of $Ra$ and $Pr$ studied in this work, as long as the flow is buoyancy-dominated, the system is in the $II_{u}$ regime with $\epsilon_{u}\sim\epsilon_{u,bulk}$. With increasing shear forcing, the shear term of (7) increases the boundary layer contribution of the kinetic dissipation due to the formation of streamwise velocity gradients close to the wall. For sufficiently strong shear, the kinetic dissipation will be dominated by the boundary layer contribution with $\epsilon_{u}\sim\epsilon_{u,BL}$ as shown in 8a. On the contrary, figure 8b shows that the thermal dissipation rate is dominated by the boundary layer contribution for the entire range of $Re_{S}$ considered in this study but the contribution reduces noticeably towards higher $Re_{S}$. For extremely strong shear, a possibility of a transition towards bulk dominance in thermal dissipation cannot be ruled out.
@@ 3.5 Friction coefficient in Rayleigh–Bénard flow
Figure 9: (a) $u_{R}^{+}$ plotted against $z^{+}$, showing the onset of a log-layer with the grey dashed line. The velocity profile for $Ra=10^{12}$, $Pr=1.0$ is indicated in magenta. (b) Diagnostic function plotted against $z^{+}$ with the inflection point and the corresponding von Kármán constant indicated. (c) $\delta^{+}$ plotted against $Re_{R}$ shows good agreement with $\delta^{+}=(1/2)Re_{R}^{1/2}$, indicated with the dashed grey line. (d) $C_{R}$ plotted against $Re_{\delta}$ showing a collapse without any prefactors. The solid grey line indicates $C_{R}\sim 4Re_{\delta}^{-2/3}$ and the dashed grey line indicates the modified Prandtl (1932) friction law given by (39). (e) $C_{R}/a$ plotted against $Re_{R}$, showing the Prandtl–Blasius scaling (23) at low $Re_{R}$ with the solid grey line and modified Prandtl (1932) friction law given by (39) at high $Re_{R}$ with the dashed grey line. (f) $C_{\epsilon}/b$ given by (40b) plotted against $Re_{R}$. The solid grey line indicates $C_{\epsilon}\sim Re_{R}^{-1/2}$ for the regime with $\epsilon_{u}\sim\epsilon_{u,BL}$, the dashed grey line indicates the modified Prandtl (1932) friction law (39). At higher $Re_{R}$, $\epsilon_{u}\sim\epsilon_{u,bulk}$, leading to $C_{\epsilon}$ becoming independent of $Re_{R}$, as indicated with the dash-dotted and dotted grey lines.
So far, we have considered only the laminar Prandtl–Blasius type kinetic boundary layers in the buoyancy-dominated regime with the boundary layer only becoming turbulent in highly shear-dominated regime. However, we now consider the possibility of turbulent kinetic boundary layers even in pure RB flow which corresponds to the so-called “ultimate” regime in RB flow (Kraichnan, 1962; Grossmann & Lohse, 2011; Roche, 2020; Lohse & Shishkina, 2023). While the existence of logarithmic temperature profiles has already been observed at high $Ra$ (Grossmann & Lohse, 2012; Ahlers et al., 2012, 2014), at the highest thermal forcing studied in this work with $Ra=10^{12}$, we start to also observe the onset of a log-layer (see figure 9a) given by
@@
Since the log-layer is not fully developed, only an estimate of the modified von Kármán constant $k$ can be obtained, namely from the inflection point of the diagnostic function plotted in figure 9b, giving $k\approx 1$. Correspondingly, the intercept $B\approx 5.2$ is found by fitting the data as shown in figure 9a.
We can now propose a relation for $C_{R}$ which is analogous to relation (33) for $C_{S}$ and (27) for $C_{T}$ by stating that in the limit of highly turbulent boundary layer, $u_{R}^{+}=U_{R}$ at $z^{+}=\delta^{+}$, where $\delta$ is the wall-normal distance to the peak velocity shown in figure 4. This gives us
@@
Although the values of $k$ and $B$ might be different, the general form of this equation suggests universality in the behaviour of wall-bounded flows. By plotting $\delta^{+}$ against $Re_{R}$, we find a very good fit with the scaling $\delta^{+}\sim Re_{R}^{1/2}$ as shown in figure 9c. We can only provide this scaling empirically because more data at extremely high $Ra$ are needed to understand the dynamics of the turbulent boundary layer which is difficult due to the high computational expense of such numerical simulations. Plotting $C_{R}$ against $Re_{\delta}$ in 9d, we find a nice collapse of the data for different aspect ratios and geometries. At low $Re_{R}$, $C_{R}$ seems to scale as $Re_{\delta}^{-2/3}$ which can be obtained from the empirical scaling $\delta^{+}\sim Re_{R}^{1/2}$ shown in figure 9c and the assumption of a Prandtl–Blasius type scaling of $C_{R}\sim Re_{R}^{-1/2}$. At high $Re_{R}$, we see good qualitative agreement between the data and the modified Prandtl (1932) friction law given by (39).

### v2

@@ 3.3 Large scale circulation
Figure 6: Visualisation of $\boldsymbol{\tau_{w}^{\prime}}$ for the CRB system with $Ra=10^{7}$ and $Pr=1$. Plots (c), (e), and (g) are in the buoyancy-dominated regime and plot (i) is in the shear-dominated regime. The colour bars indicate the magnitude of $\boldsymbol{\tau_{w}^{\prime}}$ while the black arrows in the magnified panels (b), (d), (f), (h) and (j) indicate the direction of $\boldsymbol{\tau_{w}^{\prime}}$. The black lines in (c-h) indicate the slope of $\bigl{(}Re_{w}$/$Re_{R}\bigr{)}^{-1}$. This figure is also available as an interactive JFM Notebook.
@@
Figure 7: The probability distribution $\phi(\alpha)$ of the flow orientation angle $\alpha$ for all heights $0\leq z/\lambda_{\theta}\leq Nu$ for $Ra=10^{7}$, $Pr=1$, $\Gamma_{x}=48$ and $\Gamma_{y}=24$ in (a) RB flow (i.e., with $Re_{S}=0)$, (b,d) buoyancy-dominated regime with $Re_{S}/Re_{R}\approx 2.11$, and (c,e) shear-dominated regime with $Re_{S}/Re_{R}\geq 10$. Plots (b,c) are for the CRB system while the plots (d,e) are for the PRB system. The black dashed lines indicate the height of the thermal boundary layer and the colour bar indicates the magnitude of the probability. The probability distribution $\phi(\alpha)$ at the thermal boundary layer height plotted against angle $\alpha$ for various $Ra$ and $Pr$ in (f) RB flow, (g) buoyancy-dominated regime, and (h) shear-dominated regime. The solid lines are the CRB system with $\Gamma_{x}=48$ and $\Gamma_{y}=24$, the dashed lines are for the CRB system with $\Gamma_{x}=8$ and $\Gamma_{y}=4$, and the dotted lines are for the PRB system with $\Gamma_{x}=8$ and $\Gamma_{y}=4$. This figure is also available as an interactive JFM Notebook.
@@ 3.4 Dissipation rates
Figure 8: (a) Ratio $\bigl{(}H^{2}\epsilon_{\theta}\kappa^{-1}\Delta^{2}\bigr{)}/Nu$ showing good agreement with equation 6 indicated using the solid black line, and (b) ratio $\bigl{(}H^{4}\epsilon_{u}/\nu^{3}\bigr{)}/\bigl{(}(Nu-1)RaPr^{-2}+C_{S}Re_{S}^{3}\bigr{)}$ plotted against $Re_{S}$ showing good agreement with equation 7 of the manuscript shown using the solid black line. (c) Boundary layer contribution to the global kinetic dissipation rate $\epsilon_{u,BL}$ and (d) the boundary layer contribution to the global thermal dissipation rate $\epsilon_{\theta,BL}$. The circle markers are for the CRB system, while plus markers are for the PRB system. This figure is also available as an interactive JFM Notebook.
We now investigate the bulk and boundary layer contributions to the global kinetic and thermal dissipation rates described in §2.1. We validate the rigiorous relations given by (6) and (7) using the data obtained from numerical simulations as shown in 8a and figure 8b, respectively. For the range of $Ra$ and $Pr$ studied in this work, as long as the flow is buoyancy-dominated, the system is in the $II_{u}$ regime with $\epsilon_{u}\sim\epsilon_{u,bulk}$. With increasing shear forcing, the shear term of (7) increases the boundary layer contribution of the kinetic dissipation due to the formation of streamwise velocity gradients close to the wall. For sufficiently strong shear, the kinetic dissipation will be dominated by the boundary layer contribution with $\epsilon_{u}\sim\epsilon_{u,BL}$ as shown in 8c. On the contrary, figure 8d shows that the thermal dissipation rate is dominated by the boundary layer contribution for the entire range of $Re_{S}$ considered in this study but the contribution reduces noticeably towards higher $Re_{S}$. For extremely strong shear, a possibility of a transition towards bulk dominance in thermal dissipation cannot be ruled out.
@@ 3.5 Friction coefficient in Rayleigh–Bénard flow
Figure 9: (a) $u_{R}^{+}$ plotted against $z^{+}$, showing the onset of a log-layer with the grey dashed line. The velocity profile for $Ra=10^{12}$, $Pr=1.0$ is indicated in magenta. (b) Diagnostic function plotted against $z^{+}$ with the inflection point and the corresponding von Kármán constant indicated. (c) $\delta^{+}$ plotted against $Re_{R}$ shows good agreement with $\delta^{+}=(1/2)Re_{R}^{1/2}$, indicated with the dashed grey line. (d) $C_{R}$ plotted against $Re_{\delta}$ showing a collapse without any prefactors. The solid grey line indicates $C_{R}\sim 4Re_{\delta}^{-2/3}$ and the dashed grey line indicates the modified Prandtl (1932) friction law given by (39). (e) $C_{R}/a$ plotted against $Re_{R}$, showing the Prandtl–Blasius scaling (23) at low $Re_{R}$ with the solid grey line and modified Prandtl (1932) friction law given by (39) at high $Re_{R}$ with the dashed grey line. (f) $C_{\epsilon}/b$ given by (40b) plotted against $Re_{R}$. The solid grey line indicates $C_{\epsilon}\sim Re_{R}^{-1/2}$ for the regime with $\epsilon_{u}\sim\epsilon_{u,BL}$, the dashed grey line indicates the modified Prandtl (1932) friction law (39). At higher $Re_{R}$, $\epsilon_{u}\sim\epsilon_{u,bulk}$, leading to $C_{\epsilon}$ becoming independent of $Re_{R}$, as indicated with the dash-dotted and dotted grey lines. This figure is also available as an interactive JFM Notebook.
So far, we have considered only the laminar Prandtl–Blasius type kinetic boundary layers in the buoyancy-dominated regime with the boundary layer only becoming turbulent in highly shear-dominated regime. However, we now consider the possibility of turbulent kinetic boundary layers even in pure RB flow which corresponds to the so-called “ultimate” regime in RB flow (Kraichnan, 1962; Grossmann & Lohse, 2011; Roche, 2020; Lohse & Shishkina, 2023). While the existence of logarithmic temperature profiles has already been observed at high $Ra$ (Grossmann & Lohse, 2012; Ahlers et al., 2012, 2014), it is yet to be seen if logarithmic behaviour given by
@@
can be observed in the velocity profiles.
At the highest thermal forcing studied in this work with $Ra=10^{12}$, we start to also observe some hints of what could possibly be the onset of a log-layer (see figure 9a), although it cannot be conclusively confirmed with the currently available data. Assuming that such a log-layer could exist, an estimate of the modified von Kármán constant $k$ is obtained from the inflection point of the diagnostic function plotted in figure 9b, giving $k\approx 1$. Correspondingly, the intercept $B\approx 5.2$ is found by fitting the data as shown in figure 9a.
In the presence of such a logarithmic layer, we can now hypothesise about a relation for $C_{R}$ which is analogous to relation (33) for $C_{S}$ and (27) for $C_{T}$ by stating that in the limit of highly turbulent boundary layer, $u_{R}^{+}=U_{R}$ at $z^{+}=\delta^{+}$, where $\delta$ is the wall-normal distance to the peak velocity shown in figure 4. This gives us
@@
If applicable to high $Ra$ RB flow, the general form of this equation could suggest universality in the behaviour of wall-bounded flows even though the values of $k$ and $B$ might be different from those observed for pipe or channel flows. By plotting $\delta^{+}$ against $Re_{R}$, we find a very good fit with the scaling $\delta^{+}\sim Re_{R}^{1/2}$ as shown in figure 9c. At present, we can only provide this scaling empirically because more data at extremely high $Ra$ are needed to understand the dynamics of the turbulent boundary layer which is difficult due to the high computational expense of such numerical simulations. Plotting $C_{R}$ against $Re_{\delta}$ in 9d, we find a nice collapse of the data for different aspect ratios and geometries. At low $Re_{R}$, $C_{R}$ seems to scale as $Re_{\delta}^{-2/3}$ which can be obtained from the empirical scaling $\delta^{+}\sim Re_{R}^{1/2}$ shown in figure 9c and the assumption of a Prandtl–Blasius type scaling of $C_{R}\sim Re_{R}^{-1/2}$. At high $Re_{R}$, we see good qualitative agreement between the data and the modified Prandtl (1932) friction law given by (39).